OpenStudy (anonymous):

h(t)=cot t , itervals given, π/4,3π/4 finally got the equation written out, took me and while to figure it out, sorry. book answer is -4/pi but I have no idea how to set this up

OpenStudy (amistre64):

set up for?

OpenStudy (anonymous):

what does it mean "the answer is \[-\frac{4}{\pi}\]? there is no question, just a definition of a function \[h(t)=\cot(t)\]

OpenStudy (anonymous):

average rate of change

OpenStudy (anonymous):

oh that is a question!

OpenStudy (anonymous):

sorry, duh

OpenStudy (amistre64):

average rate of change:\[\frac{change. in. y}{change.in.x}\]from begining to end

OpenStudy (anonymous):

\[\frac{h(\frac{3\pi}{4})-h(\frac{\pi}{4})}{\frac{3\pi}{4}-\frac{\pi}{4}}\]

OpenStudy (amistre64):

\[\frac{cot(3pi/4)-cot(pi/4)}{3pi/4-pi/4}\]

OpenStudy (amistre64):

latex is slowing me down lol

OpenStudy (anonymous):

i can do the bottom: \[\frac{3\pi}{4}-\frac{\pi}{4}=\frac{\pi}{2}\]

OpenStudy (amistre64):

can we factor out a cot up top?

OpenStudy (anonymous):

lol, right

OpenStudy (amistre64):

tanpi/4 = 1; flip for cot = 1 tan3pi/4 = -1; flip for cot = -1

OpenStudy (amistre64):

-1-1 = -2 for the top

OpenStudy (amistre64):

\[\frac{-2}{pi/2}=-\frac{4}{pi}\]

OpenStudy (anonymous):

I swear math makes me seriously think I may be retarded, I just cant understand it. Very difficult for my 30 year old brain. LOL...

OpenStudy (amistre64):

just wait till your my age :)

OpenStudy (anonymous):

I hope I catch on fairly quickly because I have to pass this course.

OpenStudy (anonymous):

Thank you both very much!!

OpenStudy (anonymous):

30!! i never met anyone so old....

OpenStudy (anonymous):

I know! Old as dirt! LOL, at least that is how Cal. makes me feel anyway, I should know this stuff :(

OpenStudy (anonymous):

Sorry to jump in after 7months on this question, but I'm having trouble understanding how you factored out the cot for the top. I'm getting stuck at \[\cot(3\pi/4)−\cot(\pi/4)\] How did you arrive at 1 and -1 for the top? I plugged it into my calculator and it came up something like 0.01something.

OpenStudy (amistre64):

i believe that 7 months ago that was a joke :) pi/4 radians is equivalent to 45 degrees. the tangent of an angle is the slope of the line that the radius makes with the positive x axis. the cotangent is just the inverse of the tangent.  tan(pi/4) = n/n = 1, the inverse of 1 is: 1 when we have three pi/4 s, we end up with a 135 degree angle;  as such; tan(3pi/4) = -n/n = -1, the inverse is: -1 -1 - (1) = -2